# Adversarial Score Identity Distillation: Rapidly Surpassing the Teacher in One Step

### v1

@@ Abstract
Score identity Distillation (SiD) is a data-free method that has achieved state-of-the-art performance in image generation by leveraging only a pretrained diffusion model, without requiring any training data. However, the ultimate performance of SiD is constrained by the accuracy with which the pretrained model captures the true data scores at different stages of the diffusion process. In this paper, we introduce SiDA (SiD with Adversarial Loss), which not only enhances generation quality but also improves distillation efficiency by incorporating real images and adversarial loss. SiDA utilizes the encoder from the generator’s score network as a discriminator, boosting its ability to distinguish between real images and those generated by SiD. The adversarial loss is batch-normalized within each GPU and then combined with the original SiD loss. This integration effectively incorporates the average “fakeness” per GPU batch into the pixel-based SiD loss, enabling SiDA to distill a single-step generator either from scratch or by fine-tuning an existing one. SiDA converges significantly faster than its predecessor when trained from scratch, and swiftly improves upon the original model’s performance after an initial warmup period during fine-tuning from a pre-distilled SiD generator. This one-step adversarial distillation method has set new benchmarks for generation performance when distilling EDM diffusion models pretrained on CIFAR-10 (32x32) and ImageNet (64x64), achieving FID scores of 1.499 on CIFAR-10 unconditional, 1.396 on CIFAR-10 conditional, and 1.110 on ImageNet 64x64. Our open-source code will be integrated into the SiD codebase on GitHub.
@@ 1 Introduction
In pursuit of transcending traditional challenges associated with JSD and model-based KL divergence—which are often difficult to optimize and prone to mode-seeking behaviors, respectively—Zhou et al. (2024) have pioneered a model-based Fisher divergence within the diffusion distillation framework. This novel approach, notably data-free, has demonstrated for the first time the potential to match or even surpass the performance of teacher models in a single generation step. However, SiD and other data-free methods are based on the assumption that the score produced by teacher networks can well represent the data score. This assumption can potentially create a performance bottleneck for distilled single-step generators, especially if the teacher diffusion model is not well-trained or has limited capacity. In this paper, we aim to further enhance SiD by integrating real-world data. This integration seeks to compensate for the teacher model’s limitations in accurately representing the true score, thereby producing even more realistic generative outcomes. This enhancement broadens the practical applications of diffusion distillation in real-world scenarios.
@@ 2 Related Work
Numerous techniques have been proposed to accelerate diffusion models (Zheng et al., 2022; Lyu et al., 2022; Luhman & Luhman, 2021; Salimans & Ho, 2022; Zheng et al., 2023; Meng et al., 2023; Kim et al., 2023). Recently, one-step and few-step diffusion distillation methods have gained prominence, achieving the speed of GANs while retaining the robust capabilities of diffusion models. These methods can be broadly classified into two categories: data-free (Luo et al., 2023b; Nguyen & Tran, 2024; Zhou et al., 2024) and those requiring real images or teacher-synthesized noise-image pairs (Liu et al., 2023; Song & Dhariwal, 2023; Luo et al., 2023a; Kim et al., 2023; Sauer et al., 2023; Xu et al., 2023; Yin et al., 2024b). The latter frequently incorporate GAN-based adversarial learning as a critical component to enhance the generation performance of baseline distillation methods, which vary from consistency models (Song et al., 2023) to score distillation sampling (Poole et al., 2023; Wang et al., 2023c). Without this adversarial component, performance often degrades significantly (Kim et al., 2023; Sauer et al., 2023; Yin et al., 2024a).
@@ 4 SIDA: Adversarial Score identity Distillation
SiD employs an alternating optimization strategy, updating $\psi$ according to (5) and $\theta$ according to (6), using only generator-synthesized fake data, drawn as specified in (1), for parameter updates. This approach ensures that both components are iteratively refined with outputs generated by the model itself, without relying on real data inputs.
@@ 5 Experimental Results
Datasets. To comprehensively evaluate the effectiveness of SiDA, we utilize four representative datasets of varying scales and resolutions, as discussed in Karras et al. (2022). These datasets include CIFAR-10 ($32\times 32$; both conditional and unconditional) (Krizhevsky et al., 2009), ImageNet ($64\times 64$) (Deng et al., 2009), FFHQ ($64\times 64$) (Karras et al., 2019), and AFHQ-v2 ($64\times 64$) (Choi et al., 2020). These datasets allow us to assess SiDA’s performance across different content types and complexities, ensuring a robust evaluation of its generative capabilities.
Evaluation protocol. We assess the quality of image generation using the Fréchet Inception Distance (FID) and Inception Score (IS) (Salimans et al., 2016). Consistent with the methodology of Karras et al. (2019; 2022), FID scores are calculated using 50,000 generated samples, with the training dataset used by the EDM teacher model [^1] https://github.com/NVlabs/edm serving as the reference. Additionally, we evaluate SiD on ImageNet $64\times 64$ using the Precision and Recall metrics (Kynkäänniemi et al., 2019), where both metrics are calculated using a predefined reference batch [^2] https://openaipublic.blob.core.windows.net/diffusion/jul-2021/ref_batches/imagenet/64/VIRTUAL_imagenet64_labeled.npz [^3] https://github.com/openai/guided-diffusion/tree/main/evaluations (Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Song et al., 2023; Song & Dhariwal, 2023).
Implementation details. We implement SiDA based on the codebases of EDM (Karras et al., 2022) and SiD (Zhou et al., 2024). The score estimation network, $f_{\psi}$, is initialized using the same architecture and parameters as the pre-trained teacher score network, $f_{\phi}$, from EDM (Karras et al., 2022). For the initialization of the student generator, $G_{\theta}$, we explore two settings: one where $G_{\theta}$ is initialized with the same parameters as the pre-trained EDM teacher score network, and another where it is initialized using the distilled generator from SiD, referred to as SiD-SiDA (SiD2A).
@@
We conducted an ablation study to evaluate the impact of the gradient bias correction weight factor $\alpha$, a pivotal hyperparameter for SiD, on the performance of SiDA. Specifically, we tested a spectrum of $\alpha$ values $[-0.25,0.0,0.5,0.75,1.0,1.2,1.5]$ during the distillation of the EDM model pretrained on CIFAR-10 in an unconditional context. Figure 2 demonstrates how the model’s performance varies with different $\alpha$ values over the first 20k iterations at a batch size of 256, assessed using FID and IS. We determined that an $\alpha$ value of 1.0 yielded the best performance, prompting us to adopt this setting for subsequent experiments of SiDA over various datasets. In contrast, lower $\alpha$ values such as $-0.25$ or 0 resulted in suboptimal performance, while high values like 1.2 or 1.5 led to training instability and divergence. Interestingly, the performance disparity for $\alpha$ values between 0.5 and 1.0 was significantly narrower than those observed in SiD, as depicted in Figure 8 of Zhou et al. (2024). These findings imply that although the adversarial loss may introduce some instability, it could also compensate for the inadequate correction of gradient biases associated with smaller $\alpha$ values.
@@
Figure 3: Evolution of FIDs for the SiD generator during the distillation of the EDM teacher model pretrained on CIFAR-10 in unconditional (left) and conditional (right) context, using $\alpha=1.0$ or $\alpha=1.2$ and a batch size of 256. The performance of EDM (35 steps), along with DMD and Diff-Instruct, is depicted with horizontal lines in purple, green, and red, respectively.
@@
Figure 4: Analogous plot to Figure 3 for ImageNet 64x64 with batch 8192 (left) , FFHQ-64x64 with batch size of 512 (middle), and AFHQ-v2-64x64 with batch size of 512 (right).
Table 4: Analogous to Table 2 for FFHQ 64x64.
Table 5: Analogous to Table 2 for AFHQ-v2 64x64.
5.2 Benchmark Performance
We begin by assessing the performance of SiDA and SiD2A, focusing on both generation quality and convergence speed across the datasets previously mentioned. Our comprehensive evaluation compares SiDA and SiD2A against top-performing deep generative models, including GANs, traditional diffusion models, and their distilled counterparts, in terms of generation quality. Qualitatively, random images generated by SiD2A in a single step are displayed from Figures 5 to 9 in the Appendix, illustrating the high quality achieved.
@@ 5.2 Benchmark Performance
It’s important to mention that SiDA’s closest competitor, GDD-I (Zheng & Yang, 2024), utilized the LPIPS (Zhang et al., 2018) encoder as the discriminator, which could potentially enhance FID performance through accidental leakage of ImageNet features. This concern, extensively discussed in Kynkäänniemi et al. (2023) and also raised by Song & Dhariwal (2023), does not apply to SiD, SiDA, or SiD2A, ensuring our results reflect genuine model capabilities.
SiDA also demonstrates accelerated convergence, as evidenced in Figures 3 and 4. The logarithmic scale plots show SiDA’s FID decreasing more rapidly than SiD’s, an efficiency that also holds when scaled up to ImageNet 64x64. Overall, SiDA consistently outperforms SiD in terms of convergence speed by a substantial margin, marking a significant acceleration, particularly noteworthy since SiD already converges at an exponential rate. This rapid improvement in both speed and performance underscores SiDA’s effectiveness and efficiency in generating high-quality images across different datasets. These results not only highlight SiD2A’s enhanced image quality and reduced iteration needs but also the overall efficiency of SiDA and SiD2A across various datasets, confirming the advantages of these advanced distillation techniques.
@@ 6 Conclusion
In this work, we introduce Score identity Distillation with Adversarial Loss (SiDA), a novel distillation-based framework designed to enhance the efficiency and effectiveness of diffusion model distillation processes. By synergizing the objectives of both Score identity Distillation (SiD) and Diffusion-GAN, SiDA enables high-quality image generation while substantially reducing the number of iterations required by traditional score distillation methods. Through extensive experimentation on diverse datasets, including CIFAR-10, ImageNet, FFHQ, and AFHQ-v2, SiDA has consistently outperformed both baseline diffusion models and their distilled counterparts. This superiority is evidenced by improved Fréchet Inception Distance (FID) and Inception Score (IS) metrics. SiDA represents a significant advancement in the field of efficient diffusion distillation, enhancing the performance of distilled models. Future research may extend to exploring additional distillation techniques and the potential application of SiDA in other generative tasks and domains.

### v2

@@ Abstract
Score identity Distillation (SiD) is a data-free method that has achieved state-of-the-art performance in image generation by leveraging only a pretrained diffusion model, without requiring any training data. However, the ultimate performance of SiD is constrained by the accuracy with which the pretrained model captures the true data scores at different stages of the diffusion process. In this paper, we introduce SiDA (SiD with Adversarial Loss), which not only enhances generation quality but also improves distillation efficiency by incorporating real images and adversarial loss. SiDA utilizes the encoder from the generator’s score network as a discriminator, boosting its ability to distinguish between real images and those generated by SiD. The adversarial loss is batch-normalized within each GPU and then combined with the original SiD loss. This integration effectively incorporates the average “fakeness” per GPU batch into the pixel-based SiD loss, enabling SiDA to distill a single-step generator either from scratch or by fine-tuning an existing one. SiDA converges significantly faster than its predecessor when trained from scratch, and swiftly improves upon the original model’s performance after an initial warmup period during fine-tuning from a pre-distilled SiD generator. This one-step adversarial distillation method establishes new benchmarks in generation performance when distilling EDM diffusion models pretrained on CIFAR-10 (32x32) and ImageNet (64x64), achieving FID scores of 1.499 on CIFAR-10 unconditional, 1.396 on CIFAR-10 conditional, and 1.110 on ImageNet 64x64. It sets record-low FID scores when distilling EDM2 models trained on ImageNet (512x512), surpassing even the largest teacher model, EDM2-XXL, which had an FID of 1.81 with classifier-free guidance. Our SiDA’s results—achieved without classifier-free guidance—record FID scores of 2.156 for EDM2-XS, 1.669 for EDM2-S, 1.488 for EDM2-M, and 1.465 for EDM2-L, demonstrating significant improvements across all model sizes. Our open-source code will be integrated into the SiD codebase.
@@ 1 Introduction
In pursuit of transcending traditional challenges associated with JSD and model-based KL divergence—which are often difficult to optimize and prone to mode-seeking behaviors, respectively—Zhou et al. (2024b) have pioneered a model-based Fisher divergence within the diffusion distillation framework. This novel approach, notably data-free, has demonstrated for the first time the potential to match or even surpass the performance of teacher models in a single generation step. However, SiD and other data-free methods are based on the assumption that the score produced by teacher networks can well represent the data score. This assumption can potentially create a performance bottleneck for distilled single-step generators, especially if the teacher diffusion model is not well-trained or has limited capacity. In this paper, we aim to further enhance SiD by integrating real-world data. This integration seeks to compensate for the teacher model’s limitations in accurately representing the true score, thereby producing even more realistic generative outcomes. This enhancement broadens the practical applications of diffusion distillation in real-world scenarios.
@@ 2 Related Work
Numerous techniques have been proposed to accelerate diffusion models (Zheng et al., 2022; Lyu et al., 2022; Luhman & Luhman, 2021; Salimans & Ho, 2022; Zheng et al., 2023; Meng et al., 2023; Kim et al., 2023). Recently, one-step and few-step diffusion distillation methods have gained prominence, achieving the speed of GANs while retaining the robust capabilities of diffusion models. These methods can be broadly classified into two categories: data-free (Luo et al., 2023b; Nguyen & Tran, 2024; Zhou et al., 2024b) and those requiring real images or teacher-synthesized noise-image pairs (Liu et al., 2023; Song & Dhariwal, 2023; Luo et al., 2023a; Kim et al., 2023; Sauer et al., 2023; Xu et al., 2023; Yin et al., 2024b). The latter frequently incorporate GAN-based adversarial learning as a critical component to enhance the generation performance of baseline distillation methods, which vary from consistency models (Song et al., 2023) to score distillation sampling (Poole et al., 2023; Wang et al., 2023c). Without this adversarial component, performance often degrades significantly (Kim et al., 2023; Sauer et al., 2023; Yin et al., 2024a).
@@ 4 SIDA: Adversarial Score identity Distillation
An alternative theoretical justification for this loss has recently been explored in Luo et al. (2024), demonstrating that the SiD loss can be derived under certain assumptions about the data generation mechanism, specifically involving stop-gradient conditions. SiD employs an alternating optimization strategy, updating $\psi$ according to (5) and $\theta$ according to (6), using only generator-synthesized fake data, drawn as specified in (1), for parameter updates. This approach ensures that both components are iteratively refined with outputs generated by the model itself, without relying on real data inputs.
@@ 5 Experimental Results
Datasets and Teacher Models. To comprehensively evaluate the effectiveness of SiDA, we first utilize four representative datasets of varying scales and resolutions and the EDM diffusion models pretrained on them, as discussed in Karras et al. (2022). These datasets include CIFAR-10 ($32\times 32$; both conditional and unconditional) (Krizhevsky et al., 2009), ImageNet ($64\times 64$) (Deng et al., 2009), FFHQ ($64\times 64$) (Karras et al., 2019), and AFHQ-v2 ($64\times 64$) (Choi et al., 2020).
Additionally, we distill EDM2 diffusion models of varying sizes that were pretrained on ImageNet ($512\times 512$), following details provided by Karras et al. (2024). We compare these distilled models to their corresponding teacher models and to a concurrent work of Lu & Song (2024) on EDM2 distillation. The models range from approximately 100 million to 1 billion parameters. By utilizing these diverse datasets and model scales, we thoroughly assess SiDA’s performance across different content types and complexities, ensuring a robust evaluation of its generative capabilities.
Evaluation protocol. We assess the quality of image generation using the Fréchet Inception Distance (FID) and Inception Score (IS) (Salimans et al., 2016). Following Karras et al. (2024), we also evaluate models trained on ImageNet ($512\times 512$) using $\text{FD}_{\text{DINOv2}}$, the Fréchet distance computed in the feature space of DINOv2 (Oquab et al., 2023). Consistent with the methodology of Karras et al. (2019; 2022; 2024), these scores are calculated using 50,000 generated samples, with the training dataset used by the EDM or EDM2 teacher model [^1] https://github.com/NVlabs/edm serving as the reference. Additionally, we evaluate SiDA on ImageNet $64\times 64$ using the Precision and Recall metrics (Kynkäänniemi et al., 2019), where both metrics are calculated using a predefined reference batch [^2] https://openaipublic.blob.core.windows.net/diffusion/jul-2021/ref_batches/imagenet/64/VIRTUAL_imagenet64_labeled.npz [^3] https://github.com/openai/guided-diffusion/tree/main/evaluations (Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Song et al., 2023; Song & Dhariwal, 2023). In line with SiD, we periodically evaluate the FID during distillation and preserve the generators with the lowest FID. To ensure accuracy, we conduct re-evaluations across 10 independent runs to obtain reliable metrics. We recommend this more statistically rigorous approach over reporting only the best metrics observed across independent runs or within the distillation process.
Implementation details. We implement SiDA using the SiD codebase (Zhou et al., 2024b), incorporating necessary functions adapted from either EDM (Karras et al., 2022) or EDM2 (Karras et al., 2024). The score estimation network, $f_{\psi}$, is initialized using the same architecture and parameters as the pre-trained teacher score network, $f_{\phi}$, from either EDM or EDM2. For the initialization of the student generator, $G_{\theta}$, we explore two settings: one where $G_{\theta}$ is initialized with the same parameters as the pre-trained EDM or EDM2 teacher score network, and another where it is initialized using the distilled generator from SiD, referred to as SiD-SiDA (SiD2A).
@@
We conducted an ablation study to evaluate the impact of the gradient bias correction weight factor $\alpha$, a pivotal hyperparameter for SiD, on the performance of SiDA. Specifically, we tested a spectrum of $\alpha$ values $[-0.25,0.0,0.5,0.75,1.0,1.2,1.5]$ during the distillation of the EDM model pretrained on CIFAR-10 in an unconditional context. Figure 2 demonstrates how the model’s performance varies with different $\alpha$ values over the first 20k iterations at a batch size of 256, assessed using FID and IS. We determined that an $\alpha$ value of 1.0 yielded the best performance, prompting us to adopt this setting for subsequent experiments of SiDA over various datasets. In contrast, lower $\alpha$ values such as $-0.25$ or 0 resulted in suboptimal performance, while high values like 1.2 or 1.5 led to training instability and divergence. Interestingly, the performance disparity for $\alpha$ values between 0.5 and 1.0 was significantly narrower than those observed in SiD, as depicted in Figure 8 of Zhou et al. (2024b). These findings imply that although the adversarial loss may introduce some instability, it could also compensate for the inadequate correction of gradient biases associated with smaller $\alpha$ values.
@@
Figure 3: Evolution of FIDs for the SiD and SiDA generator during the distillation of the EDM teacher model pretrained on CIFAR-10 unconditional (left) and conditional (middle), with a batch size of 256, and on ImageNet 64x64 (right) with a batch size of 8192, using $\alpha=1.0$ or $\alpha=1.2$. The performance of EDM (35 NFEs), along with DMD and Diff-Instruct, is depicted with horizontal lines in purple, green, and red, respectively.
@@
5.2 Benchmarking EDM Distillation
We begin by assessing the performance of SiDA and SiD2A, focusing on both generation quality and convergence speed across the datasets previously mentioned. Our comprehensive evaluation compares SiDA and SiD2A against top-performing deep generative models, including GANs, traditional diffusion models, and their distilled counterparts, in terms of generation quality. Qualitatively, random images generated by SiD2A in a single step are displayed from Figures 6 to 10 in the Appendix, illustrating the high quality achieved.
@@ 5.2 Benchmarking EDM Distillation
It’s worth noting that SiDA’s close competitor for ImageNet ($64\times 64$), GDD-I (Zheng & Yang, 2024), uses a discriminator inspired by Projected GAN (Sauer et al., 2021), integrating features from VGG16-BN and EfficientNet-lite0. Concerns have been raised, as discussed in Kynkäänniemi et al. (2023) and by Song & Dhariwal (2023), about the potential for encoders pretrained on ImageNet to inadvertently enhance FID scores through feature leakage. This potential issue does not affect SiD, SiDA, or SiD2A, ensuring our results reflect true model performance without such confounds.
SiDA also demonstrates accelerated convergence, as evidenced in Figures 3 and 5. The logarithmic scale plots show SiDA’s FID decreasing more rapidly than SiD’s, an efficiency that also holds when scaled up to ImageNet 64x64. Overall, SiDA consistently outperforms SiD in terms of convergence speed by a substantial margin, marking a significant acceleration, particularly noteworthy since SiD already converges at an exponential rate. This rapid improvement in both speed and performance underscores SiDA’s effectiveness and efficiency in generating high-quality images across different datasets. These results not only highlight SiD2A’s enhanced image quality and reduced iteration needs but also the overall efficiency of SiDA and SiD2A across various datasets, confirming the advantages of these advanced distillation techniques.
5.3 Benchmarking EDM2 Distillation
We validate the effectiveness of SiD, SiDA, and SiD2A in distilling EDM2 models of six different sizes pretrained on ImageNet 512x512. Our findings are summarized in Tables 4 and 7 and illustrated in Figure 4. Notably, SiD, our baseline single-step generator introduced six months ago on arXiv, performs comparably to the EDM2 teacher, which requires 63 NFEs, as well as to the simple, stable, and scalable consistency distillation (sCD) method of Lu & Song (2024) that uses 2 NFEs—a concurrent work to SiDA. SiD also convincingly outperforms sCD with 1 NFE, despite not yet employing CFG during distillation. Consistent with findings from EDM distillation, both SiDA and SiD2A demonstrate superior performance over SiD in distilling EDM2, achieving FIDs of 1.756 and 1.669, respectively, using the EDM2-S model with only 280M parameters. These results already surpass the EDM2-XXL model and sCD at 1.5B parameters, whose FIDs exceed 1.8.
Notably, scaling SiDA to EDM2-XL achieves a record-low FID of 1.449. While results for SiD2A on EDM2-XL and for the SiD family on EDM2-XXL are forthcoming, Figure 4 indicates the potential for further FID reductions as these results become available.
Mirroring findings from EDM distillation, SiDA also exhibits accelerated convergence when distilling EDM2, as shown in Figure 4. Logarithmic-scale plots reveal a more rapid decrease in FID for SiDA compared to SiD, an efficiency that scales effectively to ImageNet 512x512 and across generators of various sizes. Overall, SiDA consistently surpasses prior and concurrent methods in terms of both final metrics and convergence speed. This rapid improvement highlights SiDA’s effectiveness and efficiency in generating high-resolution images across various model sizes.
Table 4: FID scores and number of parameters for different methods on ImageNet (512x512).
Method CFG NFE XS (125M) S (280M) M (498M) L ( [^7] M) XL (1.1B) XXL (1.5B)
EDM2 N 63 3.53 2.56 2.25 2.06 1.96 1.91
EDM2 Y 63 2.91 2.23 2.01 1.88 1.85 1.81
sCT Y 1 - 10.13 5.84 5.15 4.33 4.29
sCT Y 2 - 9.86 5.53 4.65 3.73 3.76
sCD Y 1 - 3.07 2.75 2.55 2.40 2.28
sCD Y 2 - 2.50 2.26 2.04 1.93 1.88
SiD N 1 3.353 $\pm$ 0.041 2.707 $\pm$ 0.054 2.060 $\pm$ 0.038 1.907 $\pm$ 0.016 2.005 $\pm$ 0.024 -
SiDA N 1 2.228 $\pm$ 0.037 1.756 $\pm$ 0.023 1.546 $\pm$ 0.023 1.501 $\pm$ 0.024 1.449 $\pm$ 0.033 -
SiD2A N 1 2.156 $\pm$ 0.028 1.669 $\pm$ 0.019 1.488 $\pm$ 0.038 1.465 $\pm$ 0.028 - -
(a) EDM-XS
(b) EDM-S
(c) EDM-M
(d) EDM-L
(e) EDM-XL
(f) EDM-XXL
Figure 4: Evolution of FIDs for the SiD and SiDA generators during the distillation of the EDM2 teacher models of six different sizes pretrained on ImageNet (512x512), with a batch size of 2048 and $\alpha=1.0$. The performance of EDM2 without classifier-free guidance (CFG) and EDM2 with CFG, using 64 NFEs, along with the simple, stable, and scalable consistency distillation (sCD) method of Lu & Song (2024) with 1 and 2 NFEs, is depicted with horizontal lines in purple, red, green, and cyan respectively.
@@ 6 Conclusion
In this paper, we present Score identity Distillation with Adversarial Loss (SiDA), an innovative framework that optimizes diffusion model distillation by integrating the strengths of Score identity Distillation (SiD) and Diffusion-GAN. SiDA accelerates the generation process, significantly reducing the number of iterations required compared to conventional score distillation methods, while producing high-quality images. Extensive tests across various datasets—CIFAR-10, ImageNet (64x64 and 512x512), FFHQ, AFHQ-v2—and different scales of pretrained EDM2 models demonstrate that SiDA consistently surpasses both the base teacher diffusion models and other distilled variants. This is validated by superior Fréchet Inception Distance (FID) scores and other performance metrics. Notably, SiDA achieves these results without utilizing classifier-free guidance, suggesting potential for further enhancement with its integration. Specifically, incorporating the long-and-short guidance strategy as described by Zhou et al. (2024a) could further refine SiDA’s performance. Future work will explore expanding SiDA’s methodology to additional generative tasks and refining its capabilities through innovative distillation techniques.